Suppose the inverse supply curve for good X is Px = 120 + 0.5Qs. If a 10% ad valorem tax is placed on Good X, what is the new inverse supply curve? If a $10 excise tax is placed on Good X, what is the new inverse supply curve?


1
Expert's answer
2021-05-17T18:30:34-0400

"P_X=120+0.5Q_S"

If 10% ad-valorem tax is imposed on good x

then the new equation of the supply curve will be

"P_X-0.10P_X=120+0.5Q_S"

"0.9P_X=120+0.5Q_S"

"P_X=133.33+0.55Q_S"


If a $10 excise tax is imposed on the supply side, then the new equation of the supply curve will be

"P_X-10=120+0.5Q_S"

"P_X=130+0.5Q_S"

"P_X=130+0.5Q_S"